# Charges, complex structures, and perturbations of instantons

### v2

@@ 5.1. A divergence identity
Then the following identity holds
^$\hat{\nabla}_{a}\hat{V}^{a}={}$ $\frac{1}{6}|\hat{\nabla}\hat{\varphi}|_{\hat{g}}^{2}+|\hat{j}|_{\hat{g}}^{2}+\frac{1}{3\hat{\lambda}_{3}}\left[(\hat{\lambda}_{1}-\hat{\lambda}_{2})^{2}-2(\hat{\lambda}_{1}|\hat{A}|_{\hat{g}}^{2}+\hat{\lambda}_{2}|\hat{B}|_{\hat{g}}^{2})\right]$ (5.4)
$+\frac{2}{3}\hat{\varphi}^{AB}\hat{\varphi}^{CD}\hat{\nabla}_{AA^{\prime}}\hat{\nabla}^{A^{\prime}E}(\hat{\lambda}_{3}^{-1}\hat{\Psi}_{BCDE})$
where $\hat{A}_{a}$ and $\hat{B}_{a}$ are the 1-forms (2.5a), (2.5b) with respect to $\hat{g}_{ab}$.
@@
The spinor objects in the first line of (5.4) can be written in tensor terms by introducing the self-dual 2-form ^$\hat{\mathcal{F}}_{ab}=\sqrt{2}\,\hat{\varphi}_{AB}\hat{\epsilon}_{A^{\prime}B^{\prime}}$,
@@
Recall that the identity (5.4) is formulated in terms of the metric $\hat{g}_{ab}=\Omega^{2}g_{ab}$. To understand the last term in (5.4), we write it in terms of $g_{ab}$,
^$\hat{\nabla}^{AA^{\prime}}(\hat{\lambda}_{3}^{-1}\hat{\Psi}_{ABCD})=\hat{\lambda}_{3}^{-3}\nabla^{AA^{\prime}}\Psi_{ABCD},$ (5.6)
which implies that it vanishes if $(M,g_{ab})$ is Einstein. In the spirit of Wu [42] and LeBrun [34], it follows in this case, that in view of (3.8) and (5.6), the second line in (5.4) vanishes. Hence, integrating (5.4), taking this vanishing into account, and using the divergence theorem, we have
@@
$\hat{P}=\int_{M}\left(|\hat{j}|_{\hat{g}}^{2}+\frac{1}{3\hat{\lambda}_{3}}\left[(\hat{\lambda}_{1}-\hat{\lambda}_{2})^{2}-2(\hat{\lambda}_{1}|\hat{A}|_{\hat{g}}^{2}+\hat{\lambda}_{2}|\hat{B}|_{\hat{g}}^{2})\right]\right)\hat{\boldsymbol{\epsilon}}.$ (5.8)
Thus, if the boundary term vanishes and $\hat{P}$ is positive, then $(M,\hat{g}_{ab})$ is necessarily Kähler, since the equation |^$|\hat{\nabla}\hat{\varphi}|_{\hat{g}}^{2}=0$ implies ^$\hat{\nabla}_{a}\hat{\varphi}_{BC}=0$, so $J^{a}{}_{b}=\hat{\mathcal{F}}_{bc}\hat{g}^{ca}$ is a parallel almost-complex structure. The Einstein (or Ricci-flat) manifold $(M,g_{ab})$ is then conformally Kähler.
As mentioned, for compact Einstein manifolds with $\det(W^{+})>0$ the above result was shown by Wu [42]. Here, we followed LeBrun’s formulation [34] but dropping the compactness assumption (as in [12]) and the conditions $\nabla\cdot W^{+}=0$ and $\det(W^{+})>0$. The fact that $\det(W^{+})>0$ ensures $\hat{P}\geq 0$ follows from (2.10), since (2.10a) implies that one can choose $\hat{\lambda}_{3}>0$ and $\hat{\lambda}_{1}<0$, so, from (2.10c), ${\rm sign}\det(W^{+})=-{\rm sign}\,\hat{\lambda}_{2}$.
@@ 5.2. Perturbative expansions
If we do not assume that the boundary term in (5.7) vanishes or that $\hat{P}$ is positive, then we cannot conclude that $(M,\hat{g}_{ab})$ is Kähler. However, our goal now is to study infinitesimal Einstein deformations of (5.7) around a Hermitian instanton. We recall that our notation and treatment for perturbations was summarized in section 3.2. A perturbative expansion of the eigenvalue problem of the map $\hat{W}^{+}_{ab}{}^{cd}$ is justified by the fact that the latter is real and symmetric, see [30].
@@
Let $g_{ab}(s)$ be a smooth, one-parameter family of Einstein metrics in $M$, and consider the conformally related metrics $\hat{g}_{ab}(s)$ as in section 5.1. Suppose that $g_{ab}=g_{ab}(0)$ is Hermitian (so $\hat{g}_{ab}=\hat{g}_{ab}(0)$ is Kähler, with Kähler form ^$\hat{\mathcal{F}}_{ab}$). Then the perturbative expansion of the identities (5.4) and (5.7) around $s=0$ gives the following formulas
${\rm d}\,\delta\hat{\mathcal{F}}={}$ $0,$ (5.9)
$\frac{1}{4}\oint_{\partial M}\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\,(\delta^{2}\hat{\mathcal{F}})={}$ $\int_{M}\left[\frac{1}{24}\big{|}\delta(\hat{\nabla}\hat{\mathcal{F}})\big{|}_{\hat{g}}^{2}+\big{|}\delta\hat{j}\big{|}_{\hat{g}}^{2}\right.$
$\left.+\frac{1}{3\hat{\lambda}_{3}}\left((\delta\hat{\lambda}_{1}-\delta\hat{\lambda}_{2})^{2}+\hat{\lambda}_{3}(|\delta{\hat{A}}|_{\hat{g}}^{2}+|\delta\hat{B}|_{\hat{g}}^{2})\right)\right]\hat{\boldsymbol{\epsilon}}$ (5.10)
Furthermore, if the left hand side of (5.10) vanishes, then
@@
We first analyze the perturbative expansion of the right hand side of (5.4). Since $\hat{g}_{ab}=\hat{g}_{ab}(0)$ is Kähler, the background self-dual Weyl tensor $\hat{W}^{+}_{abcd}$ has two repeated eigenvalues: $\hat{\lambda}_{1}=\hat{\lambda}_{2}=-\hat{\lambda}_{3}/2$, where $\hat{\lambda}_{3}>0$. The parallel background Kähler form is ^$\hat{\mathcal{F}}_{ab}$. Recalling (5.5), we then have:
@@
^(s)^$\hat{\nabla}^{(s)}\hat{\mathcal{F}}(s)={}$ $\delta(\hat{\nabla}\hat{\mathcal{F}})\,s+O(s^{2}),$
@@
$\hat{A}_{a}(s)={}$ $\delta\hat{A}_{a}\,s+O(s^{2}),$
$\hat{B}_{a}(s)={}$ $\delta\hat{B}_{a}\,s+O(s^{2}).$
Replacing in (5.4), we see that the first non-trivial term is $O(s^{2})$, and is equal to the integrand on the right hand side of (5.10).
For a perturbative expansion of the left hand side of (5.4) (the divergence term), it is convenient to multiply by the volume form $\hat{\boldsymbol{\epsilon}}$, so that we can use the identity (^$(\hat{\nabla}_{a}\hat{V}^{a})\,\hat{\boldsymbol{\epsilon}}=-{\rm d}(\hat{*}\hat{V})$, where on the right side, $\hat{V}$ is interpreted as a 1-form. Using (5.5), one can also check that ^$\hat{*}\hat{V}=-\frac{1}{2}\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}}$. Thus:
(^$(\hat{\nabla}_{a}\hat{V}^{a})\,\hat{\boldsymbol{\epsilon}}=\tfrac{1}{2}{\rm d}(\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}}).$
The perturbative expansion of ^^$\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}}$ is:
(^^$(\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}})(s)=\delta(\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}})\,s+\frac{1}{2!}\delta^{2}(\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}})\,s^{2}+O(s^{3}),$
where
$\delta(\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}})={}$^$\hat{\mathcal{F}}\wedge\hat{*}{\rm d}(\delta\hat{\mathcal{F}}),$
$\delta^{2}(\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}})={}$^$\hat{\mathcal{F}}\wedge\hat{*}{\rm d}(\delta^{2}\hat{\mathcal{F}})+2\delta\hat{\mathcal{F}}\wedge\hat{*}{\rm d}(\delta\hat{\mathcal{F}})+2\hat{\mathcal{F}}\wedge(\delta\hat{*}){\rm d}(\delta\hat{\mathcal{F}}).$
Since there is no $O(s)$ term on the right side of (5.7), equation (5.9) follows. Equality of the $O(s^{2})$ terms gives (5.10).
To prove (5.11), note that every term on the right hand side of (5.10) is positive definite, so if the boundary term on the left vanishes, then every term on the right side vanishes as well. In particular $\delta(\hat{\nabla}\hat{\mathcal{F}})=0$, which is eq. (5.11).
Finally, we analyze the boundary term. In the compact case with $\lambda>0$, there is no boundary, so the left hand side of (5.10) vanishes. For the ALF case, we write the 3-form in (5.10) in indices,
[^^$[\hat{\mathcal{F}}\wedge\hat{*}{\rm d}(\delta^{2}\hat{\mathcal{F}})]_{abc}=\frac{3}{2}\hat{\mathcal{F}}_{[ab}\hat{\epsilon}_{c]klm}\hat{g}^{kd}\hat{g}^{le}\hat{g}^{mf}\partial_{d}(\delta^{2}\hat{\mathcal{F}})_{ef}.$ (5.12)
@@
Let $\alpha_{abc}$ denote the 3-form (5.12) restricted to $\Sigma_{r}$, then $\alpha_{abc}=f\mu_{abc}$ for some function $f$, where $\mu_{abc}$ is the volume 3-form on $\Sigma_{r}$ induced from the volume 4-form of $g_{ab}$. Using (5.13) we have $|\mu|_{g}=\sqrt{3!}$, hence $|f|=\frac{1}{\sqrt{3!}}|\alpha|_{g}$. To estimate $|\alpha|_{g}$, we first use that ^$\hat{\mathcal{F}}_{ab}=\Omega^{2}\mathcal{F}_{ab}$, $\hat{\epsilon}_{abcd}=\Omega^{4}\epsilon_{abcd}$, $\hat{g}^{ab}=\Omega^{-2}g^{ab}$, and that $|\mathcal{F}|_{g}=\sqrt{2}$, $|\epsilon|_{g}=\sqrt{4!}$, $|g|_{g}=2$, then
|^$|\hat{\mathcal{F}}|_{g}=\sqrt{2}\,\Omega^{2},\qquad|\hat{\epsilon}|_{g}=\sqrt{4!}\,\Omega^{4},\qquad|\hat{g}^{-1}|_{g}=2\,\Omega^{-2}.$ (5.14)
@@
Combining (5.14) and (5.15), we get ^$\hat{\mathcal{F}}_{ab}=O_{g}(r^{-2})$, $\hat{\epsilon}_{abcd}=O_{g}(r^{-4})$, and $\hat{g}^{ab}=O_{g}(r^{2})$, where we recall that the notation $t_{a_{1}...a_{n}}=O_{g}(r^{\alpha})$ means that there is a constant $C$ such that $|t|_{g}\leq Cr^{\alpha}$ for sufficiently large $r$, see [3, Def. 2.6]. Thus, the factor ^$\hat{\mathcal{F}}_{[ab}\hat{\epsilon}_{c]klm}\hat{g}^{kd}\hat{g}^{le}\hat{g}^{mf}$ in (5.12) is $O_{g}(1)$. To estimate the remaining factor $\partial_{d}(\delta^{2}\hat{\mathcal{F}})_{ef}$, we first use that $\delta^{2}\hat{\mathcal{F}}=\delta^{2}(\Omega^{2})\mathcal{F}+\Omega^{2}\delta^{2}\mathcal{F}+2\delta(\Omega^{2})\delta\mathcal{F}$, so every term on the right side is $O_{g}(r^{-2})$. In addition, due to the skew-symmetry of (5.12), $\partial_{d}$ can be replaced by the Levi-Civita connection $\nabla_{d}^{(s)}$of $g_{ab}(s)$, so we can use that by assumption each derivative has extra fall-off, hence $\nabla_{d}^{(s)}(\delta^{2}\hat{\mathcal{F}})_{ef}=O_{g}(r^{-3})$.
Summarizing, so far we have that the 3-form (5.12) on $\Sigma_{r}$ is $\alpha_{abc}\propto|\alpha|_{g}\mu_{abc}$ and $|\alpha|_{g}=O_{g}(r^{-3})$. From [10, Def. 1.1], $\mu_{abc}$ goes like $r^{2}$. Hence, the integrand in the left hand side of (5.10) is order $r^{-1}$, thus it vanishes in the limit $r\to\infty$. ∎
@@
To relate $\delta\hat{\mathcal{F}}_{ab}$ to the closed 2-form $\delta\hat{\omega}_{ab}$ of section 3.2, choose $Z_{ab}(s)$ to be an eigenform of $W^{+}_{ab}{}^{cd}(s)$ in (3.17). We get $\Sigma_{ab}(s)=-2\Psi_{2}(s)Z_{ab}(s)$, and
@@
[…]
@@
Eq. (5.11) can be written more explicitly in terms of perturbed tensor fields, by using the perturbative expansions given in section 3.2. Recalling formulas (3.14) and (3.15) (formulated in terms of $\hat{g}_{ab}$), and using ^$\hat{\nabla}_{a}\hat{\mathcal{F}}_{bc}=0$, we have
@@
Note that (5.17)=0italic-(5.17italic-)$=0$ gives again ${\rm d}\,\delta\hat{\mathcal{F}}=0$.
@@
Let $g_{ab}(s)$ be a curve in $\mathscr{E}(M)$, the space of ALF Ricci-flat metrics on $M$, with $g_{ab}(0)=g_{ab}$, which is by assumption Hermitian non-Kähler. The tensor field $J^{a}{}_{b}(s)=\hat{\omega}_{bc}(s)\hat{g}^{ca}(s)$ is an almost-complex structure, and Theorem 5.3 implies that
^$\hat{\nabla}_{a}^{(s)}J^{b}{}_{c}(s)=O(s^{2}),$ (5.19)

### v3

@@ 5.1. A divergence identity
Then the following identity holds:
$\hat{\nabla}_{a}\hat{V}^{a}=\mathcal{A}+\mathcal{B}$ (5.4)
where
$\mathcal{A}={}$ $|\hat{j}|^{2}+\frac{1}{6}|\hat{\nabla}\hat{\varphi}|^{2}+\frac{1}{3\hat{\lambda}_{3}}\left[(\hat{\lambda}_{1}-\hat{\lambda}_{2})^{2}-2(\hat{\lambda}_{1}|\hat{\Gamma}^{3}{}_{1}|_{\hat{g}}^{2}+\hat{\lambda}_{2}|\hat{\Gamma}^{3}{}_{2}|_{\hat{g}}^{2})\right]$ (5.5a)
$\mathcal{B}={}$ $\frac{2}{3}\hat{\varphi}^{AB}\hat{\varphi}^{CD}\hat{\nabla}_{AA^{\prime}}(\hat{\lambda}_{3}^{-3}\nabla^{A^{\prime}E}\Psi_{BCDE})$ (5.5b)
and where $\hat{\Gamma}^{3}{}_{1}$ and $\hat{\Gamma}^{3}{}_{2}$ are the 1-forms (2.5a), (2.5b) with respect to $\hat{g}_{ab}$.
@@
The spinor objects in (5.4)-(5.5a) can be written in tensor terms by introducing the self-dual 2-form $\hat{\mathcal{F}}_{ab}=\sqrt{2}\,\hat{\varphi}_{AB}\hat{\epsilon}_{A^{\prime}B^{\prime}}$,
@@
To write the term $\mathcal{B}$ in (5.5b) in terms of tensor fields, define
$\hat{P}^{abcd}=\hat{\varphi}^{(AB}\hat{\varphi}^{CD)}\hat{\epsilon}^{A^{\prime}B^{\prime}}\hat{\epsilon}^{C^{\prime}D^{\prime}},$ (5.7)
then a calculation gives
$\mathcal{B}=-\tfrac{1}{3}\hat{P}^{abcd}\hat{\nabla}_{a}(\hat{\lambda}_{3}^{-1}\nabla_{d}E_{bc}),$ (5.8)
where $E_{bc}=R_{bc}-\frac{R}{4}g_{bc}$ is the trace-free Ricci tensor of $g_{ab}$.
@@
Recall that the identity (5.4) is formulated in terms of the metric $\hat{g}_{ab}=\Omega^{2}g_{ab}$. To understand the term $\mathcal{B}$ in (5.5b), we write it in terms of $g_{ab}$,
$\hat{\nabla}^{AA^{\prime}}(\hat{\lambda}_{3}^{-1}\hat{\Psi}_{ABCD})=\hat{\lambda}_{3}^{-3}\nabla^{AA^{\prime}}\Psi_{ABCD},$ (5.9)
which implies that it vanishes if $(M,g_{ab})$ is Einstein. In the spirit of Wu [55] and LeBrun [44], it follows in this case, that in view of (3.8) and (5.9), we have $\mathcal{B}\equiv 0$. Hence, integrating (5.4), taking this vanishing into account, and using the divergence theorem, we have
@@
$\hat{P}=\int_{M}\left(|\hat{j}|_{\hat{g}}^{2}+\frac{1}{3\hat{\lambda}_{3}}\left[(\hat{\lambda}_{1}-\hat{\lambda}_{2})^{2}-2(\hat{\lambda}_{1}|\hat{\Gamma}^{3}{}_{1}|_{\hat{g}}^{2}+\hat{\lambda}_{2}|\hat{\Gamma}^{3}{}_{2}|_{\hat{g}}^{2})\right]\right)\hat{\boldsymbol{\epsilon}}$. (5.11)
Thus, if the boundary term vanishes and $\hat{P}$ is positive, then $(M,\hat{g}_{ab})$ is necessarily Kähler, since the equation $|\hat{\nabla}\hat{\varphi}|_{\hat{g}}^{2}=0$ implies $\hat{\nabla}_{a}\hat{\varphi}_{BC}=0$, so $J^{a}{}_{b}=\hat{\mathcal{F}}_{bc}\hat{g}^{ca}$ is a parallel almost-complex structure. The Einstein (or Ricci-flat) manifold $(M,g_{ab})$ is then conformally Kähler.
As mentioned, for compact Einstein manifolds with $\det(W^{+})>0$ the above result was shown by Wu [55]. Here, we followed LeBrun's formulation [44] but dropping the compactness assumption (as in [16]) and the conditions $\nabla\cdot W^{+}=0$ and $\det(W^{+})>0$. The fact that $\det(W^{+})>0$ ensures $\hat{P}\geq 0$ follows from (2.10), since (2.10a) implies that one can choose $\hat{\lambda}_{3}>0$ and $\hat{\lambda}_{1}<0$, so, from (2.10c), ${\rm sign}\det(W^{+})=-{\rm sign}\,\hat{\lambda}_{2}$.
@@ 5.2. Perturbative expansions
Our goal now is to study infinitesimal Einstein deformations of the generic identity (5.4) around a Hermitian instanton. We recall that our notation and treatment for perturbations was summarized in section 3.2. A perturbative expansion of the eigenvalue problem of the map $\hat{W}^{+}_{ab}{}^{cd}$ is justified by the fact that the latter is real and symmetric, see [39].
@@
Let $g_{ab}(s)$ be a smooth curve of Riemannian metrics in $M$, and consider the conformally related metrics $\hat{g}_{ab}(s)$ as in section 5.1. Suppose that $g_{ab}=g_{ab}(0)$ is Hermitian-Einstein (thus $\hat{g}_{ab}=\hat{g}_{ab}(0)$ is Kähler, with Kähler form $\hat{\mathcal{F}}_{ab}$), and that $\frac{{\rm d}}{{\rm d}s}g_{ab}(s)|_{s=0}$ is a linearized Einstein deformation, that is
$\left.E_{ab}[g(s)]\right|_{s=0}=0,\qquad\left.\frac{{\rm d}}{{\rm d}s}E_{ab}[g(s)]\right|_{s=0}=0,$ (5.12)
where $E_{ab}[g(s)]=R_{ab}[g(s)]-\frac{1}{4}R[g(s)]\,g_{ab}(s)$ is the trace-free Ricci tensor of $g_{ab}(s)$. Then, integrating the identity (5.4) and taking a perturbative expansion around $s=0$, we get the following formulas:
$\int_{M}\hat{\nabla}_{a}\hat{V}^{a}=\left\{\oint_{\partial M}\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\delta\hat{\mathcal{F}}\right\}s+\left\{\oint_{\partial M}\delta^{2}(\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\,\hat{\mathcal{F}})\right\}\frac{s^{2}}{2!}+O(s^{3}),$ (5.13a)
$\int_{M}\mathcal{A}=\int_{M}\left\{\frac{|\delta(\hat{\nabla}\hat{\mathcal{F}})|^{2}}{24}+|\delta\hat{j}|^{2}+\frac{(\delta\hat{\lambda}_{1}-\delta\hat{\lambda}_{2})^{2}}{3\hat{\lambda}_{3}}+\frac{(|\delta\hat{\Gamma}^{3}{}_{1}|^{2}+|\delta\hat{\Gamma}^{3}{}_{2}|^{2})}{3}\right\}\frac{s^{2}}{2!}+O(s^{3})$ (5.13b)
$\int_{M}\mathcal{B}=\left\{-\frac{1}{3}\oint_{\partial M}(\hat{\lambda}_{3}^{-1}\hat{P}^{abcd}\nabla_{d}\delta^{2}E_{bc}){\rm d}\Sigma_{a}\right\}\frac{s^{2}}{2!}+O(s^{3}).$ (5.13c)
In particular, it follows that
${\rm d}\,\delta\hat{\mathcal{F}}=0.$ (5.14)
Furthermore, if the boundary terms of order $O(s^{2})$ in (5.13a) and (5.13c) vanish, then
@@
[…]
The perturbative expansion of $\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}}$ is:
$(\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}})(s)=\delta(\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}})\,s+\delta^{2}(\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}})\,\frac{s^{2}}{2!}+O(s^{3}),$ (5.16)
where
[…]
This gives (5.13a).
We now focus on (5.13b). Since $\hat{g}_{ab}=\hat{g}_{ab}(0)$ is Kähler, the background self-dual Weyl tensor $\hat{W}^{+}_{abcd}$ has a double eigenvalue: $\hat{\lambda}_{1}=\hat{\lambda}_{2}=-\hat{\lambda}_{3}/2$, where $\hat{\lambda}_{3}>0$. The parallel background Kähler form is $\hat{\mathcal{F}}_{ab}$. Recalling (5.6), we then have:
@@
$\hat{\nabla}^{(s)}\hat{\mathcal{F}}(s)={}$ $\delta(\hat{\nabla}\hat{\mathcal{F}})\,s+O(s^{2}),$
@@
[…]
Replacing in (5.5a), we see that the first non-trivial term is $O(s^{2})$, and is equal to the integrand on the right hand side of (5.13b).
We continue with (5.13c). We shall use the expression (5.8), i.e.
$\mathcal{B}(s)=-\tfrac{1}{3}\widehat{P}^{acbd}(s)\widehat{\nabla}^{(s)}_{a}\left[\widehat{\lambda}_{3}(s)^{-1}\nabla^{(s)}_{d}E_{bc}(s)\right]$
where we are abbreviating $E_{bc}(s)=E_{bc}[g(s)]$. Using (5.12), we have $\mathcal{B}(s)=(\delta^{2}\mathcal{B})\,\frac{s^{2}}{2!}+O(s^{3})$, and
$\delta^{2}\mathcal{B}={}$ $\left.\frac{{\rm d}^{2}}{{\rm d}s^{2}}\mathcal{B}(s)\right|_{s=0}$
$={}$ $-\tfrac{1}{3}\hat{P}^{acbd}(0)\hat{\nabla}^{(0)}_{a}\left[\hat{\lambda}_{3}(0)^{-1}\nabla^{(0)}_{d}\delta^{2}E_{bc}\right]$
$={}$ $-\tfrac{1}{3}\hat{\nabla}^{(0)}_{a}\left[\hat{\lambda}_{3}(0)^{-1}\hat{P}^{acbd}(0)\nabla^{(0)}_{d}\delta^{2}E_{bc}\right]$
where in the second line we used (5.12) and $\delta^{2}E_{bc}=\frac{{\rm d}^{2}}{{\rm d}s^{2}}E_{bc}(s)\big|_{s=0}$, and in the third line we used that $\hat{\nabla}^{(0)}_{a}\hat{P}^{bcde}(0)=0$. Integrating and using the divergence theorem, formula (5.13c) follows.
To show (5.14), we use the identity $\int_{M}(\hat{\nabla}_{a}\hat{V}^{a})(s)=\int_{M}\mathcal{A}(s)+\int_{M}\mathcal{B}(s)$ (recall (5.4)), together with (5.13). Since $\int_{M}\mathcal{A}(s)$ and $\int_{M}\mathcal{B}(s)$ do not have $O(s)$ terms, equation (5.14) follows.
To prove (5.15), suppose that the $O(s^{2})$ boundary terms coming from $\int_{M}(\hat{\nabla}_{a}\hat{V}^{a})(s)$ and $\int_{M}\mathcal{B}(s)$ vanish. Then, since every term in the $O(s^{2})$ part of the right hand side of (5.13b) is positive definite, every such term must vanish as well. In particular $\delta(\hat{\nabla}\hat{\mathcal{F}})=0$, which is eq. (5.15).
Finally, we now prove that, for compact and ALF instantons, the $O(s^{2})$ boundary terms of $\int_{M}(\hat{\nabla}_{a}\hat{V}^{a})(s)$ and $\int_{M}\mathcal{B}(s)$ vanish.
In the compact case with $\lambda>0$, there is no boundary, so the boundary terms vanish automatically.
[…]
$[\hat{\mathcal{F}}\wedge\hat{*}{\rm d}(\delta^{2}\hat{\mathcal{F}})]_{abc}=\frac{3}{2}\hat{\mathcal{F}}_{[ab}\hat{\epsilon}_{c]klm}\hat{g}^{kd}\hat{g}^{le}\hat{g}^{mf}\partial_{d}(\delta^{2}\hat{\mathcal{F}})_{ef}.$ (5.18)
@@
Let $\alpha_{abc}$ denote the 3-form (5.18) restricted to $\Sigma_{r}$, then $\alpha_{abc}=f\mu_{abc}$ for some function $f$, where $\mu_{abc}$ is the volume 3-form on $\Sigma_{r}$ induced from the volume 4-form of $g_{ab}$. Using (5.19) we have $|\mu|_{g}=\sqrt{3!}$, hence $|f|=\frac{1}{\sqrt{3!}}|\alpha|_{g}$. To estimate $|\alpha|_{g}$, we first use that $\hat{\mathcal{F}}_{ab}=\Omega^{2}\mathcal{F}_{ab}$, $\hat{\epsilon}_{abcd}=\Omega^{4}\epsilon_{abcd}$, $\hat{g}^{ab}=\Omega^{-2}g^{ab}$, and that $|\mathcal{F}|_{g}=\sqrt{2}$, $|\epsilon|_{g}=\sqrt{4!}$, $|g|_{g}=2$, then
$|\hat{\mathcal{F}}|_{g}=\sqrt{2}\,\Omega^{2},\qquad|\hat{\epsilon}|_{g}=\sqrt{4!}\,\Omega^{4},\qquad|\hat{g}^{-1}|_{g}=2\,\Omega^{-2}.$ (5.20)
@@
Combining (5.20) and (5.21), we get $\hat{\mathcal{F}}_{ab}=O_{g}(r^{-2})$, $\hat{\epsilon}_{abcd}=O_{g}(r^{-4})$, and $\hat{g}^{ab}=O_{g}(r^{2})$, where we recall that the notation $t_{a_{1}...a_{n}}=O_{g}(r^{\alpha})$ means that there is a constant $C$ such that $|t|_{g}\leq Cr^{\alpha}$ for sufficiently large $r$, see [4, Def. 2.6]. Thus, the factor $\hat{\mathcal{F}}_{[ab}\hat{\epsilon}_{c]klm}\hat{g}^{kd}\hat{g}^{le}\hat{g}^{mf}$ in (5.18) is $O_{g}(1)$. To estimate the remaining factor $\partial_{d}(\delta^{2}\hat{\mathcal{F}})_{ef}$, we first use that $\delta^{2}\hat{\mathcal{F}}=\delta^{2}(\Omega^{2})\mathcal{F}+\Omega^{2}\delta^{2}\mathcal{F}+2\delta(\Omega^{2})\delta\mathcal{F}$, so every term on the right side is $O_{g}(r^{-2})$. In addition, due to the skew-symmetry of (5.18), $\partial_{d}$ can be replaced by the Levi-Civita connection $\nabla_{d}^{(s)}$of $g_{ab}(s)$, so we can use that by assumption each derivative has extra fall-off, hence $\nabla_{d}^{(s)}(\delta^{2}\hat{\mathcal{F}})_{ef}=O_{g}(r^{-3})$.
Summarizing, so far we have that the 3-form (5.18) on $\Sigma_{r}$ is $\alpha_{abc}\propto|\alpha|_{g}\mu_{abc}$ and $|\alpha|_{g}=O_{g}(r^{-3})$. From [14, Def. 1.1], $\mu_{abc}$ goes like $r^{2}$. Hence, the integrand in the $O(s^{2})$ boundary term in (5.13a) is order $r^{-1}$, thus it vanishes in the limit $r\to\infty$.
The last step is then to analyze the $O(s^{2})$ boundary term in $\int_{M}\mathcal{B}(s)$, eq. (5.13c). To do this, we define
$\hat{U}^{a}=-\tfrac{1}{3}\hat{\lambda}_{3}^{-1}\hat{P}^{acbd}\nabla_{d}\delta^{2}E_{bc}$
so that (5.13c) becomes
$\int_{M}\mathcal{B}(s)=-\frac{s^{2}}{2!}\oint_{\partial M}\hat{*}\hat{U}+O(s^{3}).$ (5.22)
We have
$(\hat{*}\hat{U})_{abc}=-\frac{1}{3}\hat{\lambda}_{3}^{-1}\hat{\epsilon}_{abcd}\hat{g}^{dk}\hat{g}^{el}\hat{g}^{fm}\hat{g}^{gn}\hat{P}_{klmn}\nabla_{e}\delta^{2}E_{fg}$ (5.23)
Using that the ALF assumption gives $\hat{\lambda}_{3}^{-1}=O_{g}(r)$, $\hat{\epsilon}_{abcd}=O_{g}(r^{-4})$, $\hat{g}^{dk}=O_{g}(r^{2})$, $\hat{P}_{klmn}=O_{g}(r^{-4})$, we find that the term $\hat{\lambda}_{3}^{-1}\hat{\epsilon}_{abcd}\hat{g}^{dk}\hat{g}^{el}\hat{g}^{fm}\hat{g}^{gn}\hat{P}_{klmn}$ in (5.23) is $O_{g}(r)$. To estimate the remaining factor $\nabla_{e}\delta^{2}E_{fg}$, we first recall the discussion around (3.14), which gives
$\nabla_{e}\delta^{2}E_{fg}=\nabla^{(s)}_{e}\delta^{2}E_{fg}+O(s).$
Therefore, we can now use that $\delta^{2}E_{fg}=O_{g}(r^{-3})$ and $\nabla^{(s)}_{e}\delta^{2}E_{fg}=O_{g}(r^{-4})$, where we used that by assumption each $\nabla^{(s)}_{a}-$derivative has extra fall-off. Recalling (5.23), we then get $(\hat{*}\hat{U})_{abc}=O_{g}(r^{-3})$.
Taking now a 3-surface $\Sigma_{r}=\{r=const.\}$, we have $\oint_{\partial M}\hat{*}\hat{U}=\lim_{r\to\infty}\oint_{\Sigma_{r}}\hat{*}\hat{U}$. Since, as argued before, the induced volume form is $\mu_{abc}=O_{g}(r^{2})$, then $\oint_{\Sigma_{r}}\hat{*}\hat{U}$ is order $r^{-1}$. Thus, taking the limit $r\to\infty$ we get $\oint_{\partial M}\hat{*}\hat{U}=0$.
In summary, we showed that, for ALF vacuum perturbations of ALF instantons, the $O(s^{2})$ boundary terms of $\int_{M}(\hat{\nabla}_{a}\hat{V}^{a})(s)$ and $\int_{M}\mathcal{B}(s)$ vanish. Hence, (5.15) holds for ALF instantons. ∎
@@
[…]
@@
$={}$ $\sqrt{2}\,\delta\left[\Psi_{2}(s)^{2/3}Z_{ab}(s)\right]$
$={}$ $\delta\hat{\mathcal{F}}_{ab},$ (5.24)
@@
Eq. (5.15) can be written more explicitly in terms of perturbed tensor fields, by using the perturbative expansions given in section 3.2. Recalling formulas (3.14) and (3.15) (formulated in terms of $\hat{g}_{ab}$), and using $\hat{\nabla}_{a}\hat{\mathcal{F}}_{bc}=0$, we have
@@
[…]
@@
Let $g_{ab}(s)$ be a curve of Riemannian ALF metrics on $M$, where $g_{ab}(0)=g_{ab}$ is Ricci-flat and Hermitian non-Kähler, and $\delta g_{ab}=\frac{{\rm d}}{{\rm d}{s}}g_{ab}(s)|_{s=0}$ is an infinitesimal Einstein deformation. The tensor field $J^{a}{}_{b}(s)=\hat{\omega}_{bc}(s)\hat{g}^{ca}(s)$ is an almost-complex structure, and Theorem 5.3 implies that
$\hat{\nabla}_{a}^{(s)}J^{b}{}_{c}(s)=O(s^{2}),$ (5.27)